# Impacts of aspect ratio on task accuracy in parallel coordinates

### v1

@@ IV-B Analysis methodology
Task A: Accuracy was measured as a distance (the distance between the estimated and real correlation level or binary outcome (if the estimated correlation level was within 1 category of the real level). Task B: To account for error of visual estimation in the response, responses were considered correct if they were within 2 (0-50 scale) of the correct value.
@@
Generalised linear mixed models (GLMMs) were used for analysis of the relationship between the AR, dataset size and accuracy. For Task A, where accuracy is a measure of distance between the real and estimated correlation value, an ordered beta GLMM [35] (logit link) was used. This provides directionality of response and as such can indicate whether responses are likely to be under- or over- estimated. For Task B (accuracy as binary outcome), a binomial GLMM (logit link) was used. Models were fitted with interaction between AR, sample size and Fisher-z (Task A only), with random effects for participant and question order. GLMMs were evaluated using AIC [36] and residuals were checked using DHARMa [37]. Predicted responses were generated with the ‘ggpredict’ function from the ggeffects [38] package to assess marginal effects.
V Results
V-A Task A
Fig. 3 show distinct patterns for each AR, indicating that there is a high likelihood (sample size 40: $p$ 2.7 × 10-12; 160: $p$ 9.8 x 10-5) of a real difference in proportion of correct responses dependent on AR and providing support for H1.1.
The predicted marginal effects from an ordered beta model are displayed in fig. 4. The patterns in the figure support the above results and show clear evidence of interaction between AR and sample size. ARs of 1 and 2 appear more performant at sample size 160, with responses substantially closer to the real value, particularly for correlation values greater than zero.
@@ V-A Task A
Fig. 3 and fig. 4 show strongly asymmetric responses across correlations dependent on AR, supporting H1.2. Fig. 4 shows the ordering of ARs at each correlation level is largely consistent, but variance is not consistent across correlation levels (Fisher-z). Figure 4 indicates responses are slight under-estimates for both negative and positive correlations (i.e. correlation is estimated as being slightly less strong than the real value). This effect is asymmetric and there is clear divergence in response by AR - as AR decreases, responses are more likely to be substantially under-estimated for positive correlations. However, the effect varies by sample size. For a sample size of 40, responses are likely to be closer to the real value for ARs $>$1 than for ARs of 1 or lower. Notably there is a clear falloff in accuracy for AR $\leq$ 1 where correlation is greater than zero. This provides partial support for H1.3. At sample size 160, responses for AR 4 indicate large over-estimates of positive correlation, and notably zero correlation tends to be estimated as positive. The drop in performance for AR 1 at correlation $>$ 0 is much less evident, and there appears to be little evidence for a difference in performance for AR $\geq$ 1 at the highest levels of correlation, with a substantial overlap in confidence intervals.
Overlap of confidence intervals at the most negative correlations indicate that there is unlikely to be a substantial real difference between ARs, with only the extremes (AR 0.25 and 4) showing a likely difference in performance. This is consistent across sample sizes.
@@
Participant is a substantial random effect (variance logOR 0.71), with the pooled model likely to be insufficient (Chisq $p$ 2.2e-16). Trial number (i.e. question order within factors) is insignificant as a random effect.
Figure 3: Task A: Real and estimated Fisher-z scores by AR (rows) and sample size (cols) for AR 0.25, 1 and 4. See supplementary material for plot of full range of ARs evaluated.
Figure 4: Task A: Predicted and actual mean distance between estimated and actual correlation
V-B Task B
@@
Friedman tests for number of correct responses by AR at each sample size (sample size 10: $p$ 3.4 x 10-15; 20: $p$ 1.1 x 10-13; 40: $p$ 1.5 x 10-12) provide strong evidence that AR has an effect on successful tracing of values across axes (H2.1).
@@
Fig. 5 shows marginal probability of an accurate response by AR and sample size. For all sample sizes, AR 1 is most accurate. There is strong evidence for a substantial reduction in accuracy at AR 0.25 and 4 (supporting H2.2), with some limited evidence for a reduction in accuracy at AR 0.5 and 1. The expected probability of accurate responses does appear to interact with both AR and sample size. There is good evidence for a substantial reduction in accuracy at AR 2 compared to AR 1, but no apparent reduction at AR 0.5 for a sample size of 40. This asymmetry is not apparent for sample sizes 10 or 20.
@@
Figure 5: Task B: Predicted and actual probability correct by log AR and sample size (number of polylines)
VI Discussion
@@
The results for Task A indicate that correlation perception is likely to be substantially influenced by the AR of the plot. Further, the number of polylines displayed (sample size) interacts with the AR - for low numbers of polylines an AR of $\geq$ 2 is much more likely to lead to accurate correlation estimation than an AR of 1. This difference appears to be mitigated by a larger sample size, with closely matched performance across AR 1, 2 and 4. This provides some support for the hypothesis that smaller between-line angles and length differences make it easier to estimate ‘closeness-to-parallel’ and reduces salience of line crossings, but where there are many more polylines the larger pool of potential comparisons reduces the difference. For the smallest AR (0.25), the interaction with sample size appears much less significant, and plots at AR 0.25 are notably poor for estimating anything other than strong negative correlations, again appearing to support the increased angle-difference and line-length hypothesis unless the pattern formed is particularly distinctive, as is the case with the negative correlation ‘X’ pattern. Indeed, as Harrison et al. [1] show, negative correlations are perceived much more strongly than positive in a PCP, so it is possible this effect counterbalances the reduced salience of the ‘X’ pattern for high AR.
VI-A Asymmetric responses
@@
Fig. 3 and fig. 4 indicate that where there is no correlation, responses for AR $\leq$ 1 and AR 4 tend to skew negative or positive respectively, with a notably strong effect for AR 4 at a sample size of 160. While Type I errors of this kind are problematic, in exploratory analysis tasks (where the goal is identifying potentially interesting relationships) this may not present a significant downside as the effects are relatively weak and correlation could be confirmed by additional indicators. Further, for AR 0.25, responses to positive correlations are clustered around 0 (fig. 3), with fig. 4 showing a substantial under-estimation of correlation strength for AR 0.25 and 0.5, indicating that for such small horizontal distances and higher between-line angles the pattern formed by positive correlations is hard to estimate.
@@
Much less asymmetry in response is shown for value tracing tasks, where performance is maximised at AR between 0.5 and 1. Decline in performance is most strongly apparent for high AR (4), particularly for sample size of 40. This fits with the hypothesis that the separation between axes makes lines hard to follow and it is plausible that the visual clutter introduced at the higher sample size makes the falloff in accuracy more acute. Potentially the reduced between-line angle may also lead to less readily apparent separation between lines with similar values. Similarly, for very low AR (0.25) the decline in accuracy is plausibly due to the difficulty of separating individual lines where the horizontal distance is very small.
@@
Further work will be required to understand why higher sample sizes do not see a reduction (and potentially a small increase) in accuracy at AR 0.5 compared to 1, which is not apparent in the smaller sample sizes. Although this evaluation used static images, it is possible that even with brushing techniques to highlight the polyline of interest, some reduction in accuracy of rapid value estimation is possible for large AR ($>$4) due to separation distance of axes.
VI-B Limitations

### v2

@@ 4.2 Analysis methodology
Task A: Accuracy was measured as the distance between the estimated and real correlation level. For example, a true Fisher-z value of -1.5 with a response of -1 (’–’ in 2a) produces an accuracy measure of 0.5 - i.e. the participant slightly underestimated the absolute magnitude of the strongly negative correlation. Similarly, for a true Fisher-z value of 0.5 with a response of 1 (’++’ in 2a) produces an accuracy measure of 0.5 - i.e. the participant slightly overestimated the absolute magnitude of the weakly positive correlation. For the simple test of correct responses, we also convert this to a binary outcome (if the estimated correlation level was within 1 category of the real level). Task B: Responses were measured as a binary outcome (correct or incorrect). To account for error of visual estimation in the response, responses were considered correct if they were within 2 (0-50 scale) of the correct value. Note that a continuous error scale (rather than binary correct/incorrect) cannot be used here as the most likely source of error is tracing an incorrect line, and any resultant error will depend on the endpoint of that line (which may be at any point on the scale).
@@
Generalised linear mixed models (GLMMs) were used for analysis of the relationship between the aspect ratio, dataset size and accuracy. These allow the inclusion of random effects to account for the repeated measures within participants. For Task A, where accuracy is a measure of distance between the real and estimated correlation value, an ordered beta GLMM [36] (logit link) was used. This is the most appropriate fit for the distribution of the response variable, which is on a bounded scale. The model enables evaluation of direction and magnitude of response and as such can indicate whether responses are likely to be under- or over- estimated. For Task B (accuracy as binary outcome), a binomial GLMM (logit link) was used. Models were fitted with interaction between aspect ratio, sample size (both tasks) and Fisher-z (Task A only) - we expect that sample size will affect perception of the differing patterns formed by positive or negative correlations and that the effect of aspect ratio will be dependent on sample size and correlation. To account for variation in performance between participants, a random effect for participant (intercept only) was included. Further, given the use of the same dataset for all levels of aspect ratio, we include a trial number variable as a random effect to account for potential learning effects of question order. See table 1 for model details.
Table 1: Analysis models
Task Model
Task A $response\_accuracy\sim fisher\_z*aspect\_ratio*sample\_size+(1|participant\_id)+(1|trial\_num)$
Task B $response\_correct\sim aspect\_ratio*sample\_size+(1|participant\_id)+(1|trial\_num)$
Fitted models for analysis of Task A (top) and Task B (bottom), using the R notation (see [37]).
GLMMs were evaluated using AIC [38] for relative comparisons of models with and without interactions, and residuals were checked using DHARMa [39]. Models with complex higher-order interactions can be difficult to interpret directly, as coefficients in non-linear models with interactions do not map straightforwardly to changes on the response scale. We therefore follow standard practice [40] by evaluating marginal predictions at the population level (fixed effects only, with random effects held at their expected value of zero), representing the expected response for an average participant in the study. We generate adjusted predicted responses for each level of the independent variables (e.g. sample size = 40 or 160, aspect ratio = 0.25, 0.5, etc, Fisher-z = -1.5, -1, etc) with the ‘ggpredict’ [41] function from the ggeffects [42] package. This provides population-level predictions from the full model ensuring that interaction effects are included in the estimates.
To evaluate the hypothesised interaction between aspect ratio and accuracy (H1.2), we formalise inferences from the fitted model for Task A with a measure comparing the variance of predicted responses. At each sample size (40 and 160), we fit a simple linear model to the predicted response data for each Fisher-z score, weighting by the inverse of the variance of the predicted value (link scale). Taking each pair of Fisher-z values, we then compare the variance of the residuals with an F-test to indicate whether the differences in variances are due to chance i.e. does the ’spread’ of results by aspect ratio differ dependent on the true Fisher-z?
5 Results
5.1 Task A
H1.1: Aspect ratio has an effect on accurate estimation of correlation in a PCP.
Fig. 3 shows response percentage distributions for true Fisher-z by Fisher-z as estimated by participants. Values on the diagonal are correct, with values below and right of the diagonal being under-estimation and values above and left being over-estimation. There are distinct patterns for each aspect ratio, indicating that there is a high likelihood of a real difference in proportion of correct responses dependent on aspect ratio and providing support for H1.1. Skillings-Mack test results for sample size 40 stat. 59.9 $p$ 3.1 $\times$ 10-12 and 160: stat. 23.6 $p$ 9.6 x 10-5 indicate that these differences are unlikely to be due to chance.
Marginal predictions from the model are displayed in fig. 4 (see supplementary for table of values). These represent population-level predictions evaluated at specific factor levels (e.g., sample size = 40 or 160, aspect ratio, Fisher-z), with other covariates held at representative values. Confidence intervals indicate uncertainty in the estimation of the fixed effects. This indicates the expected population mean error in response by sample size group at each aspect ratio and level of Fisher-z score tested. Observed (unadjusted) sample means computed directly from the raw (post-QC) responses (no random-effects adjustment) are included as a sanity check to compare with population-level model predictions.
The error between estimated and actual correlation (figure 4) indicates direction of mean expected response and is dependent on the real correlation, with negative errors indicating overestimation for negative correlations and underestimation for positive ones. For example, for a response distance of 0.5 a Fisher-z of -1 ($r$ of -0.762) is perceived as a Fisher-z of -0.5 ($r$ of -0.462); a Fisher-z of 1 ($r$ of 0.762) is perceived as 1.5 ($r$ of 0.905).
@@ 5.1 Task A
As such, figure 4 indicates responses are slight under-estimates for both negative and positive correlations (i.e. correlation is estimated as being slightly less strong than the real value). This effect is asymmetric and there is clear divergence in response by aspect ratio - as aspect ratio decreases, responses are more likely to be substantially under-estimated for positive correlations. However, the effect varies by sample size.
For a real Fisher-z of 1, the error of estimation decreases as aspect ratio increases - for sample size 40, AR 0.25 is likely to lead to underestimating the true value by over 1 - users estimate Fisher-z values of 1 ($r$ of 0.762) around 0 (-0.09). At AR 4 however, the expected mean underestimation is much smaller at 0.22 - so users estimate values of 1 at 0.78 ($r$ of 0.653).
H1.2: Aspect ratio interacts with correlation to affect accuracy of estimated correlation.
Fig. 3 and fig. 4 show the ordering of aspect ratios at each correlation level is largely consistent, but variance is not consistent across correlation levels (Fisher-z).
For each real Fisher-z score (x-axis) the responses have varying bounds (for example, at the lowest Fisher-z, the error in responses can only ever indicate a Fisher-z closer to zero, with a positive distance to the true value). We therefore consider the spread of expected error (distance to true value) by aspect ratio at each level of real Fisher-z. This spread appears to increase with Fisher-z (with the exception of Fisher-z 1.5 at sample size 160), suggesting that aspect ratio interacts with Fisher-z and supporting H1.2. Considering the top plot in figure 4 (sample size of 40), for a true Fisher-z of -1.5, the confidence intervals of the predicted mean are overlapping for all aspect ratios and so there is no evidence for a real difference in mean, and this is consistent across both sample sizes. Conversely, at a true Fisher-z of 1 ($r$ of 0.762) there is little overlap of confidence intervals across aspect ratios, providing strong evidence for a real difference in the true means.
This gradual divergence of expected means and non-overlapping CIs as the Fisher-z increases reflects the Skillings-Mack test results. To further evaluate the interaction between aspect ratio and correlation estimation accuracy, we conducted F-tests of pairwise variance differences of the predicted response error for each real Fisher-z value (see section 4 for details). Table 2 presents the results of these F-tests, highlighting significant differences in variance across different levels of Fisher-z. For a sample size of 40, there is strong evidence that differences between strongly negative and strongly positive correlations are unlikely to be due to chance. The same effect appears also to be present, though more weakly so, for a sample size of 160.
Evidence of interaction between aspect ratio and sample size is apparent in this differing response variability in figure 4.
H1.3: estimated correlation will more closely match the real value in PCPs with a high aspect ratio (>>1) than in aspect ratios of 1 or lower.
For a sample size of 40, responses appear likely to be closer estimates of the real value when aspect ratio exceeds 1. Specifically, aspect ratios of 2 and 4 show reduced estimation error, particularly for positive correlations. Further, there is a clear falloff in accuracy for AR $\leq$ 1 where correlation is greater than zero. This provides partial support for H1.3. Aspect ratios 0.25 and 0.5 have substantial under-estimation of positive correlations, but the drop in performance for AR 1 at correlation >> 0 is much less evident. There appears to be little evidence for a difference in performance for AR $\geq$ 1 at the highest correlation (Fisher-z of 1.5), with a substantial overlap in confidence intervals (AR 1 -0.33 CI95 -0.43 to -0.24; AR 2 -0.27 CI95 -0.40 to -0.13; AR 4 -0.21 CI95 -0.34 to -0.08). Responses for AR 4 indicate large overestimates of positive correlation, and notably zero correlation tends to be estimated as positive (0.42 CI95 0.29 - 0.55).
@@
Participant is a substantial random effect (variance logOR 0.71), with the pooled model likely to be insufficient (Chisq $p$ 2.2 x 10-16). Trial number (i.e. question order within factors) is insignificant as a random effect.
Table 2: Task A: Pairwise F-tests of predicted variance by sample size and real Fisher-z. Direction indicates which of the Fisher-z 1 or Fisher-z 2 values were greater. Sig. column indicates P-values under the thresholds of $5\times 10^{-2}$ (*), $5\times 10^{-3}$ (**), $5\times 10^{-4}$ (***).
Fisher-z 1 Fisher-z 2 F stat P-value Sig. Direction
Sample size: 40
-1.5 -1 $2.7$ $1.8\times 10^{-1}$ V1 >> V2
-1.5 -0.5 $5.9$ $5.6\times 10^{-2}$ V2 >> V1
-1.5 0 $4.3$ $9.3\times 10^{-2}$ V2 >> V1
-1.5 0.5 $13.8$ $1.3\times 10^{-2}$ * V2 >> V1
-1.5 1 $21.8$ $5.6\times 10^{-3}$ * V2 >> V1
-1.5 1.5 $16.4$ $9.6\times 10^{-3}$ * V2 >> V1
-1 -0.5 $15.8$ $1.0\times 10^{-2}$ * V2 >> V1
-1 0 $11.5$ $1.8\times 10^{-2}$ * V2 >> V1
-1 0.5 $36.8$ $2.1\times 10^{-3}$ ** V2 >> V1
-1 1 $58.0$ $8.5\times 10^{-4}$ ** V2 >> V1
-1 1.5 $43.6$ $1.5\times 10^{-3}$ ** V2 >> V1
-0.5 0 $1.4$ $3.8\times 10^{-1}$ V1 >> V2
-0.5 0.5 $2.3$ $2.2\times 10^{-1}$ V2 >> V1
-0.5 1 $3.7$ $1.2\times 10^{-1}$ V2 >> V1
-0.5 1.5 $2.8$ $1.7\times 10^{-1}$ V2 >> V1
0 0.5 $3.2$ $1.4\times 10^{-1}$ V2 >> V1
0 1 $5.1$ $7.3\times 10^{-2}$ V2 >> V1
0 1.5 $3.8$ $1.1\times 10^{-1}$ V2 >> V1
0.5 1 $1.6$ $3.3\times 10^{-1}$ V2 >> V1
0.5 1.5 $1.2$ $4.4\times 10^{-1}$ V2 >> V1
1 1.5 $1.3$ $3.9\times 10^{-1}$ V1 >> V2
Sample size: 160
-1.5 -1 $1.7$ $3.1\times 10^{-1}$ V2 >> V1
-1.5 -0.5 $2.8$ $1.7\times 10^{-1}$ V2 >> V1
-1.5 0 $7.5$ $3.9\times 10^{-2}$ * V2 >> V1
-1.5 0.5 $12.5$ $1.6\times 10^{-2}$ * V2 >> V1
-1.5 1 $10.2$ $2.3\times 10^{-2}$ * V2 >> V1
-1.5 1.5 $3.7$ $1.2\times 10^{-1}$ V2 >> V1
-1 -0.5 $1.6$ $3.2\times 10^{-1}$ V2 >> V1
-1 0 $4.3$ $9.3\times 10^{-2}$ V2 >> V1
-1 0.5 $7.2$ $4.1\times 10^{-2}$ * V2 >> V1
-1 1 $5.9$ $5.7\times 10^{-2}$ V2 >> V1
-1 1.5 $2.2$ $2.4\times 10^{-1}$ V2 >> V1
-0.5 0 $2.6$ $1.9\times 10^{-1}$ V2 >> V1
-0.5 0.5 $4.4$ $9.0\times 10^{-2}$ V2 >> V1
-0.5 1 $3.6$ $1.2\times 10^{-1}$ V2 >> V1
-0.5 1.5 $1.3$ $4.0\times 10^{-1}$ V2 >> V1
0 0.5 $1.7$ $3.2\times 10^{-1}$ V2 >> V1
0 1 $1.4$ $3.9\times 10^{-1}$ V2 >> V1
0 1.5 $2.0$ $2.6\times 10^{-1}$ V1 >> V2
0.5 1 $1.2$ $4.2\times 10^{-1}$ V1 >> V2
0.5 1.5 $3.4$ $1.3\times 10^{-1}$ V1 >> V2
1 1.5 $2.7$ $1.8\times 10^{-1}$ V1 >> V2
Figure 3: Task A: Real (true) and estimated Fisher-z scores by aspect ratio (facet rows and colour) and sample size (40 and 160, facet columns). Each facet shows a combination of aspect ratio and sample size, and a facet then includes the true Fisher-z (x-axis) and participant estimated Fisher-z (y-axis). Bubble area indicates proportion of responses at each estimated Fisher-z value for the specific true Fisher-z value.
Figure 4: Task A: Marginal (population-level) model predictions of signed error (y-axis, Fisher-z) by true Fisher-z (x-axis) across aspect ratios (colour) and sample sizes (top facet sample size 40, bottom 160), with 95% CIs, plotted together with the corresponding observed (unadjusted) mean errors. For any given true Fisher-z value (x-axis), the distance of the error (y-axis) indicates the magnitude of the mean error in estimated value by participants. A positive value indicates over-estimation, a negative under-estimation.
5.2 Task B
H2.1: Aspect ratio has an effect on successful tracing of values across multiple axes.
@@ 5.2 Task B
Friedman tests for number of correct responses by aspect ratio at each sample size (sample size 10: $p$ 3.4 x 10-15; 20: $p$ 1.1 x 10-13; 40: $p$ 1.5 x 10-12) provide strong evidence that aspect ratio has an effect on successful tracing of values across axes (H2.1).
H2.2: as aspect ratio decreases below 1 or increases above 1 successful value tracing will become less likely.
@@
Fig. 5 shows marginal probability of an accurate response by aspect ratio and sample size for the population (see supplementary for table of values). Predictions are marginal, evaluated at specific factor levels (e.g., sample size = 10, 40 or 160, aspect ratio), with other covariates held at representative values. For all sample sizes, AR 1 is most accurate. There is strong evidence for a substantial reduction in accuracy at AR 0.25 (size 10 OR 0.17; size 20 OR 0.09; size 40 OR 0.22) and 4 (size 10 OR 0.13; size 20 OR 0.39; size 40 OR 0.1) which are unlikely to be due to chance (see table 3), supporting H2.2. There is some very limited evidence for a reduction in accuracy at AR 0.5 and 1 for sample size 10 (OR 0.55). The expected probability of accurate responses does appear to interact with both aspect ratio and sample size. There is good evidence for a substantial reduction in accuracy at AR 2 compared to AR 1 for sample size 40 (OR 0.28, $p$ 5.3 x 10-4), but no apparent reduction at AR 0.5 (OR 1.08). This asymmetry is not apparent for sample sizes 10 or 20.
Table 3: Task B: Pairwise evaluation of estimated accuracy by aspect ratio. Sig. column indicates P-values under the thresholds of $5\times 10^{-2}$ (*), $5\times 10^{-3}$ (**), $5\times 10^{-4}$ (***).
Sample size AR 1 AR 2 P-value Sig.
10 0.25 0.5 $1.5\times 10^{-3}$ **
10 0.25 1 $1.5\times 10^{-6}$ ***
10 0.25 2 $1.6\times 10^{-4}$ ***
10 0.25 4 $4.2\times 10^{-1}$
10 0.5 1 $1.4\times 10^{-1}$
10 0.5 2 $5.9\times 10^{-1}$
10 0.5 4 $3.3\times 10^{-4}$ ***
10 1 2 $4.0\times 10^{-1}$
10 1 4 $4.5\times 10^{-6}$ ***
10 2 4 $6.3\times 10^{-5}$ ***
20 0.25 0.5 $6.1\times 10^{-8}$ ***
20 0.25 1 $3.4\times 10^{-10}$ ***
20 0.25 2 $6.1\times 10^{-8}$ ***
20 0.25 4 $1.2\times 10^{-4}$ ***
20 0.5 1 $3.8\times 10^{-1}$
20 0.5 2 $1.0\times 10^{0}$
20 0.5 4 $1.1\times 10^{-1}$
20 1 2 $3.9\times 10^{-1}$
20 1 4 $1.5\times 10^{-2}$ *
20 2 4 $1.1\times 10^{-1}$
40 0.25 0.5 $1.7\times 10^{-5}$ ***
40 0.25 1 $4.3\times 10^{-5}$ ***
40 0.25 2 $5.4\times 10^{-1}$
40 0.25 4 $2.1\times 10^{-2}$ *
40 0.5 1 $8.3\times 10^{-1}$
40 0.5 2 $2.6\times 10^{-4}$ ***
40 0.5 4 $5.9\times 10^{-11}$ ***
40 1 2 $5.3\times 10^{-4}$ **
40 1 4 $1.9\times 10^{-10}$ ***
40 2 4 $3.2\times 10^{-3}$ **
T-test of differences in estimated predictors by sample size
@@
Figure 5: Task B: Marginal effects - model-based marginal predictions and ’real’ (observed, unadjusted) probability correct by aspect ratio and sample size (number of polylines)
6 Discussion
@@
The results for Task A indicate that correlation perception is likely to be substantially influenced by the aspect ratio of the plot. The effects of this can be large - for example, for an aspect ratio of 1, underestimating a true correlation of 1 ($r$ 0.762) by approximately 0.75 leads to a perceived correlation around 0.25 ($r$ 0.245). Further, the number of polylines displayed (sample size) interacts with the aspect ratio - for low numbers of polylines an aspect ratio of $\geq$ 2 is much more likely to lead to accurate correlation estimation than an aspect ratio of 1. This difference appears to be mitigated by a larger sample size, with closely matched performance across AR 1, 2 and 4. This provides some support for the hypothesis that smaller between-line angles and length differences make it easier to estimate ‘closeness-to-parallel’ and reduces salience of line crossings, but where there are many more polylines the larger pool of potential comparisons reduces the difference. For the smallest aspect ratio (0.25), the interaction with sample size appears to have a much smaller effect, and plots at AR 0.25 are notably poor for estimating anything other than strong negative correlations, again appearing to support the increased angle-difference and line-length hypothesis unless the pattern formed is particularly distinctive, as is the case with the negative correlation ‘X’ pattern. Indeed, as Harrison et al. [1] show, negative correlations are perceived much more strongly than positive in a PCP, so it is possible this effect counterbalances the reduced salience of the ‘X’ pattern for high aspect ratio.
6.1 Asymmetric responses
@@
Positive correlations are harder to accurately perceive at low aspect ratios. For AR 0.25, responses to positive correlations are clustered around 0 (fig. 3), with fig. 4 showing a substantial under-estimation of correlation strength for AR 0.25 and 0.5, indicating that for such small horizontal distances and higher between-line angles the pattern formed by positive correlations is hard to estimate. This effect is substantial, and is likely to lead to users completely missing all but the strongest positive correlations at low aspect ratios. Where there is no correlation, responses for AR $\leq$ 1 and AR 4 tend to skew negative or positive respectively, with a notably strong effect for AR 4 at a sample size of 160 (see fig. 4 and fig. 3 - right column, central value of Fisher-z 0). While Type I errors of this kind are problematic, in exploratory analysis tasks (where the goal is identifying potentially interesting relationships) this may not present a significant downside and correlation could be confirmed by additional indicators.
@@
Much less asymmetry in response is shown for value tracing tasks, where performance is maximised at aspect ratio between 0.5 and 1. Decline in performance is most strongly apparent for high aspect ratio (4), particularly for sample size of 40. This fits with the hypothesis that the separation between axes makes lines hard to follow and it is plausible that the visual clutter introduced at the higher sample size makes the falloff in accuracy more acute. Potentially the reduced between-line angle may also lead to less readily apparent separation between lines with similar values. Similarly, for very low aspect ratio (0.25) the decline in accuracy is plausibly due to the difficulty of separating individual lines where the horizontal distance is very small.
@@
Further work will be required to understand why higher sample sizes do not see a reduction (and potentially a small increase) in accuracy at AR 0.5 compared to 1, which is not apparent in the smaller sample sizes. Although this evaluation used static images, it is possible that even with brushing techniques to highlight the polyline of interest, some reduction in accuracy of rapid value estimation is possible for large aspect ratio (>>4) due to separation distance of axes.
6.2 Limitations